OpenStudy (kanwal32):

At room temperature ammonia gas at one atmospheric pressure and hydrogen chloride at P atmospheric pressure are allowed to effuse through identical pinholes from opposite ends of glass tube of one metre length and of unifrom cross section. AmmoniumChloride is first formed at a distance of 60cm from the end through which HCL gas is sent in. What is the value of P?

OpenStudy (kanwal32):

@nincompoop help

OpenStudy (kanwal32):

@Somy help

OpenStudy (kanwal32):

\[\frac{ r _{1} }{ r _{2} }=\frac{ p_1 }{ p_2 }\sqrt{\frac{ M_2 }{ M_1 }}\]

OpenStudy (somy):

idk sorry

OpenStudy (kanwal32):

\[\frac{r_1}{r_2}=\frac{ \sqrt{d_2} }{ \sqrt{d_1} }\]

OpenStudy (kanwal32):

i am using ths formula bot not getting the answer

OpenStudy (somy):

i don't even know these formulas T_T sorry

OpenStudy (kanwal32):

ok

OpenStudy (kanwal32):

do u know anyone who's knows these formulas and is good in chem

OpenStudy (kanwal32):

@nincompoop pls help

OpenStudy (kanwal32):

@dan815 can u pls help

OpenStudy (kanwal32):

@Abmon98 pls help

OpenStudy (kanwal32):

its ok @arithmetic555

OpenStudy (kanwal32):

@Abmon98 do u know the answer the answer is given as 2.2atm but my answer is different

OpenStudy (kanwal32):

@abhisar help @wio

OpenStudy (kanwal32):

@shamil98 hlp

OpenStudy (anonymous):

Break down your problem, You have effusion so you know you will be using Grahams law. Ammonia gas at room temp and one atmosphere pressure: Room temperature = 25 degrees Celsius. One Atmospheric Pressure = 1 ATM HCL pressure = P = ? Tube is 1 meter length = 100 centimeter AmmoniaChloride is formed 60cm away from the HCL. Image we can create.